OpenStudy (anonymous):

What is the solution set of 2x^2 + x = 15? a) {3, -5/2} b) {-3, 5/2} c) {2, -3/2}

OpenStudy (anonymous):

b) 2(-3)^2-3=15 ( 9x2)-3=15 18-3=15 15=15

OpenStudy (anonymous):

2x^2+x-15=0 (2x-5)(x+3)=0. x=-3, x=5/2

OpenStudy (anonymous):

2x^2+x-15=0 2x^2+6x-5x-15=0 2x(x+3)-5(x+3) (2x-5)(x+3)=0 x=-3,5/2